A 0.301 T magnetic field in a mass spectrometer causes an isotope of sodium to move in a circular path with a radius of 0.07 m. The charge on an ion is 1.60218
Answer #1

13 2 00
Answer #2

e/m = v/Br;

e*B*r/v = m;

m = 4.23* 10^-25 kg
Answer #3

The same question was resolved before with different figures. Please input your figure to get your answer. Let me know if you face any issue. Please rate if I succeeded in helping you.

old question

A 0.361 T magnetic field in a mass spectrometer causes an isotope of sodium to move in a circular path with a radius of 0.077m. The charge on an ion is 1.60218x10^-19C. If the ions are moving with a speed of 10500m/s, what is the isotope's mass? Answer in units of kg.   

ans

e/m = v/Br;

e*B*r/v = m;

m = 4.23* 10^-25 kg
Answer #4

The only force that can make the isotope move in a circular path is the magnetic force, which in this case must be acting perpendicular to the direction of motion at any instant, since the particle is moving in a circle with constant speed.

So equating centripetal force to the magnetic force,

(m*v*v)/r = q*v*B

where m is mass of the isotope, v is its speed, r is the radius in which it is moving, q is its charge, B is the magnetic field

which gives m = 1.596x10^(-25) kg